2 1 6 2. Evaluate 25x4 + + X -1.5 vx xt 3. Find f (x) given that f(x)= x°+2-e ** at (0,-1).
Answer #1

da 2 I= 1.5 + 6 |(25*** I = (25x – 2 + x x/ +60-4_1.5) da -ati = 25 26 - alna + ata + 6* - 4.50 (-4+1) +C le. I = 25x5 - a do